40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 10 16:19:03 2016, ericq, Update, SUS, ETMX watch 
    Reply  Mon Jun 13 10:01:58 2016, Steve, Update, SUS, ETMX jumps noDamping24hrs.pngETMXjumps.png
    Reply  Tue Jun 14 15:55:17 2016, ericq, Update, SUS, ETMX watch ETMX_wander.png
       Reply  Tue Jun 14 16:53:16 2016, rana, Update, SUS, ETMX watch 
       Reply  Thu Jun 16 11:10:17 2016, Steve, Update, SUS, spare ETMX optics preparation to be hanged SOSsus_Page_1.pngSOSsus_Page_2.pngETMXspareTowerFace.jpgETMXspareTowerBack.jpgmagnets-dumbbells.jpg
          Reply  Tue Jun 28 09:14:27 2016, Steve, Update, SUS, ETMX optics preparation to be rehanged nylonFCss.jpg
             Reply  Wed Jun 29 08:20:04 2016, Steve, Update, SUS, First Contact shelf life 
                Reply  Wed Jun 29 14:32:22 2016, Steve, Update, SUS, how to remove epoxy from optics 
                   Reply  Wed Jun 29 16:02:29 2016, Steve, Update, SUS, ETMX beam baffle IMG_1608.JPG
                   Reply  Thu Jun 30 10:03:16 2016, Steve, Update, SUS, how to remove epoxy from optics SOSsoakDish.jpgSOSsoakDish_tv.jpg
                      Reply  Thu Jun 30 14:31:02 2016, Chemistry, Update, SUS, NO mad-scientist.jpg
                      Reply  Tue Jul 5 09:25:31 2016, Steve, Update, SUS, how to remove epoxy from optics SOSsoakingDish.jpgSOSsoakingD.jpg
          Reply  Thu Jul 7 09:13:54 2016, Steve, Update, SUS, SOS gluing fixture SOSgluingFixRarmWend.jpgSOSgluingFixRarm.jpgSOSgluingFixLarm45.jpgSOSgluingFixLarm.jpgSOSgluingFixture.jpg
Message ID: 12229     Entry time: Wed Jun 29 16:02:29 2016     In reply to: 12227
Author: Steve 
Type: Update 
Category: SUS 
Subject: ETMX beam baffle 

In the attached photo from 2012, one can see the installed black glass baffle. According to the drawings (LIGO-DNNNXXX) this one has a clear aperture of 40 mm.

In (someplace ?) we have clean baffles with a 50 mm aperture which can be installed during this vent. In order to be more conservative, let us choose to swap these out for all 4 test masses during the upcoming vent using the green laser as an alignment guide, as Koji described at today's lunch meeting.

They are located at the top of E1 drawer cabinet

Attachment 1: IMG_1608.JPG  715 kB  Uploaded Wed Jun 29 20:28:47 2016  | Hide | Hide all
IMG_1608.JPG
ELOG V3.1.3-